A river has a current of 5 mph. A boat can make a trip of 20 miles upstream in the same time it can make a trip of 30 miles downstream find the rate of the boat in still water
Answer by jorel1380(3719)

30/x+5=20/x-5
20(x+5)=30(x-5)
20x+100=30x-150
10x=250
x=25
The speed of the boat in still water is 25 mph...